The dimerization of butadiene, , to form 1,5-cyclooctadiene is a second-order process that occurs when the diene is heated. In an experiment, a sample of 0.0083 mol of was heated in a 1.0-L flask. After 660. seconds, 15% of the butadiene had dimerized. Calculate the rate constant for this reaction. Rate constant =
Answer #1

we have:

[A]o = 0.0083 M

15% has dimerized. So, remaining is 85 %

[A] = 85 % of 0.0083 M = 0.007055 M

t = 660.0 s

use integrated rate law for 2nd order reaction

1/[A] = 1/[A]o + k*t

1/(0.007055) = 1/(0.0083) + k* 660.0

1.417*10^2 = 1.205*10^2 +k*6.6*10^2

k*6.6*10^2 = 21.26

k = 3.221*10^-2 M-1.s-1

Answer: 3.22*10^-2 M-1.s-1